OpenStudy (anonymous):

an investor invested a total of $1,900 in two mutual funds. one fund earned a 9% profit while the other earned a 5%profit. if the investors total profit was $143, how much was invested in each mutual fund?

OpenStudy (campbell_st):

you need to solve 2 questions let the amounts invested in each fund be x and y then x + y = 1900 the profits are 0.09x + 0.05y = 143 so you need to solve the equations x + y = 1900 (1) 0.09x + 0.05y = 143 (2) multiply equation 2 by 20 and subtract from equation 1 which will eliminate y x + y = 1900 - 1.8x + y = 2860 --------------------- -0.8x = 960 solve for x. when you have x substitute it into equation 1 to find y. this will give the amounts invested in each fund

OpenStudy (campbell_st):

oops should read -0.8 = -960 then solve for x